The switch is pinging as expected, but some servers are still porting down at the moment. Our hardware, network and datacenter teams are still investigating the source of the anomaly following the replacement that was carried out.
We'll keep you informed as soon as we have more information.
Posted Oct 02, 2023 - 14:31 CEST
Monitoring
Investigations revealed that the fault was due to an electrical problem, and the necessary equipment was replaced by the Datacenter team. Everything is now back to normal.
If you no longer have access to your server (via ssh, for example), you need to power off and power on to regain full access.
Posted Oct 02, 2023 - 14:06 CEST
Investigating
We have detected a switch down in PAR1. Servers in that rack currently have no public network access and are unreachable.
Posted Oct 02, 2023 - 12:48 CEST
This incident affected: Elements - Products (Elastic Metal).